can anyone help with this problem. fertilizer is to be spread at a rate of 2 pounds for 600 square feet. how many pounds of fertilizer will be needed for 900 square feet.
Answer by TimothyLamb(4379)

---
2/600 = w/900
w = (2/600)*900
w = 3 pounds
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php